40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4938     Entry time: Tue Jul 5 13:35:56 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

The PRM sus damping was restored. It's side rms motion came down from 35 to 4 mV immediately.      Lab   air quality is back to normal.

 

ELOG V3.1.3-